OpenStudy (anonymous):

Select all statements below which are true for all invertible n×n matrices A and B A. AB=BA B. (A+B)(A−B)=A2−B2 C. A+A−1 is invertible D. A3B8 is invertible E. (ABA−1)4=AB4A−1 F. (A+A−1)9=A9+A−9

OpenStudy (zarkon):

where are stuck...what have you tried?

OpenStudy (anonymous):

I have computed each statement, but I am still not getting the right combination

OpenStudy (zarkon):

have you figured out the answer to any of these?

OpenStudy (anonymous):

no, I do not understand the rules of invertible matrices

OpenStudy (zarkon):

do you remember the rules of (square) matrix multiplication...in particular that matrix multiplication is not commutative?

OpenStudy (zarkon):

also your notation is crazy..I believe I have figured out what you have typed...for example is (A+A−1)9=A9+A−9 \[(A+A^{-1})^9=A^9-A^{-9}\]

OpenStudy (anonymous):

Yes sorry I am new to this website

OpenStudy (anonymous):

C and D is always right. A and B are equivalent but not always correct, neither does F. As for E, i am sorry but i dont understand ur notation.

OpenStudy (zarkon):

If \(A=\left[\begin{array}{cc}0 & -1 \\ 1 & 0\\ \end{array}\right]\) then \(A^{-1}=\left[\begin{array}{cc}0 & 1 \\ -1 & 0\\ \end{array}\right]\) and thus \(A+A^{-1}=\left[\begin{array}{cc}0 & 0 \\ 0 & 0\\ \end{array}\right]\) which is not invertable